35. Your knowledge of TVM has made you quite the negotiator. Car dealers fear you. While purchasing a $25,000 car, the car dealer wants to finance it for you at 6% APR but you know you can get a 5% APR loan from your bank. If the terms of both loans are four years with monthly payments and compounding, how much less is your monthly payment as financed with the bank?

A. $11.39

B. $13.67

C. $16.48

D. $18.01

E. $20.68

36. Which of the following is an appropriate goal for the firm?

A. profit maximization

B.revenue maximization

C.tax minimization

D. market share maximization

E.shareholder wealth maximization

37. The monthly mortgage payment on your house is $750. Your mortgage lasts for 30 years and has a 6% APR compounded monthly. How much did you borrow?

  1. $118,000

  2. $119,754

  3. $120,433

  4. $123,686

  5. $125,094
Answer #1

35. If r = 6%/12 = 0.005

n = 4 * 12 = 48

PV = \frac{PMT}{r} * \left [ 1 - \frac{1}{(1+r)^{n}} \right ]

25,000 = \frac{PMT}{0.005} * \left [ 1 - \frac{1}{(1+0.005)^{48}} \right ]

25,000 = \frac{PMT}{0.005} *0.2129015889

25,000 = PMT * 42.58031778

PMT = \frac{25,000}{42.58031778}

PMT = $587.1257262374

If r = 5%/12 = 0.004166666667

PV = \frac{PMT}{r} * \left [ 1 - \frac{1}{(1+r)^{n}} \right ]

25,000 = \frac{PMT}{0.004166666667} * \left [ 1 - \frac{1}{(1+0.004166666667)^{48}} \right ]

25,000 = \frac{PMT}{0.004166666667} *0.1809289831

25,000 = PMT * 43.4229559405

PMT = \frac{25,000}{43.4229559405}

PMT = \$575.7323392322

Monthly payment will be (587.1257262374 - 575.7323392322) = 587.1257262374 - 575.7323392322 = $11.3933870052 less

Option A. $11.39 is correct

36. Which of the following is an appropriate goal for the firm?

E. Shareholder wealth maximization is the appropriate goal of the firm

Below options are incorrect:

A. profit maximization doesn't necessarily increase the wealth of shareholders

B.revenue maximization doesn't necessarily increase the wealth of shareholders because expenses could be more

C.tax minimization is not the goal of the firm

D. market share maximization this may be the goal of the marketing team, but not the goal of the firm

37. r = 6%/12 = 0.005

n = 30 * 12 = 360

PV = \frac{PMT}{r} * \left [ 1 - \frac{1}{(1+r)^{n}} \right ]

PV = \frac{750}{0.005} * \left [ 1 - \frac{1}{(1+ 0.005)^{360}} \right ]

PV =150,000 *0.833958072

PV =\$125,093.7108

The amount we borrowed = $125,094

Can you please upvote? Thank You :-)